An object has a kinetic energy of 275 J and a momentum of magnitude 25.0 kg * m/s. Find the (a) speed and (b) mass of the object.
1
Expert's answer
2020-10-25T18:31:26-0400

a) Let's first write the formulas for the momentum and kinetic energy:


"p=mv,""KE=\dfrac{1}{2}mv^2."

We can rewrite the formula for the kinetic energy as follows:


"KE=\dfrac{1}{2}mv^2=\dfrac{1}{2}pv."

From this formula we can find the speed of the object:


"v=\dfrac{2KE}{p}=\dfrac{2\cdot 275\ J}{25.0\ \dfrac{kgm}{s}}=22.0\ \dfrac{m}{s}."

b) Finally, from the first equation we can find the mass of the object:


"m=\dfrac{p}{v}=\dfrac{25.0\ \dfrac{kgm}{s}}{22.0\ \dfrac{m}{s}}=1.14\ kg."

Answer:

a) "v=22.0\ \dfrac{m}{s}."

b) "m=1.14\ kg."